# Low-Rank Adaptation for Foundation Models: A Comprehensive Review

### v1

@@ 5.9.1 Language-vision Learning
Language-Vision Model Based Adaptation. Sung et al. [203] adapted Vision-Language Model (VLM) by fine-tuning CLIP-BART for visual question answering and image captioning. Ji et al. [204] enhanced the adversarial robustness of VLMs through clustering-based LoRA for cross-modal retrieval. In addition, Zong et al. [205] proposed MoVA to efficiently route and fuse multiple vision foundation models (CLIP [15], DINOv2 [146], SAM [7]) through a coarse-to-fine adapter mechanism.
@@ 5.9.2 Language-Audio Learning
In Language-Audio Learning tasks, LoRA has emerged as a pivotal technique for addressing the fundamental challenge of modality gap between audio and text representations in foundation models. The applications primarily fall into two categories: speech recognition and audio content understanding and generation.
Speech Recognition. Fathullah et al. [214] introduced a direct audio embedding method where LoRA adapts the self-attention layers specifically for audio-text alignment, enabling large language models to process audio inputs effectively. Fathullah et al. [214] demonstrated that LLMs can perform multilingual ASR by prepending audio embeddings to text tokens, adapting only critical model parameters through LoRA while achieving competitive performance against specialized ASR systems. Yusuf et al. [215] proposed speculative speech recognition by combining an RNN-Transducer-based ASR system [216] with an audio-prefixed language model adapted using LoRA, allowing the model to generate speculative predictions. Palaskar et al. [217] further introduced FLoRA with modality-specific adapters that can be selectively engaged, providing a robust framework for multimodal integration.
Speech Content Understanding and Generation. Liu et al. [218] proposed LOAE for automated audio captioning (AAC), which combines CED-based audio encoding [219] with LLaMA for text decoding, bridged by Q-Former [220], and optimized through LoRA-based adaptation. Qin et al. [188] introduced ATFLRec, a multimodal recommender system that leverages separate LoRA modules to efficiently fine-tune audio and text modalities in LLMs, demonstrating superior performance through their fusion-based approach.
@@ 6 Challenges and Discussion
While LoRA has demonstrated remarkable success in efficiently adapting foundation models across diverse domains, several critical challenges and opportunities further further investigation.
Theoretical Understanding. Current theoretical frameworks primarily focus on simplified settings or specific architectures [72], leaving gaps in our understanding of LoRA’s behavior in more complex scenarios. For instance, the interaction between LoRA adaptations and the pretrained model’s knowledge is not fully understood, particularly in terms of how LoRA preserves useful features while modifying task-specific ones. Developing more comprehensive theoretical frameworks that can explain LoRA’s effectiveness across different architectures and tasks would be valuable for guiding future improvements.
@@
Computational Efficiency. The scalability of LoRA becomes increasingly crucial as LLMs continue to expand. Managing concurrent adaptations while handling variable-length sequences presents significant challenges in memory utilization and computational overhead. The dynamic management of adapter modules and KV cache tensors can lead to memory fragmentation and increased I/O costs [129]. Advanced serving architectures and optimization techniques are essential for maintaining low latency in production environments, particularly when dealing with multiple concurrent adaptation requests.
Robustness and Verification. The deployment of LoRA-adapted models in critical applications necessitates robust verification methods. Current research inadequately addresses model behavior under distribution shifts and adversarial conditions [223]. Developing rigorous uncertainty quantification methods and formal verification techniques becomes paramount, especially for high-stakes applications in healthcare and autonomous systems where model reliability directly impacts human safety.
@@
Future research directions should focus on: establishing comprehensive theoretical frameworks that unify various LoRA design aspects; developing automated architecture search methods for optimal adapter configurations; creating efficient serving infrastructures for large-scale deployments; implementing standardized evaluation protocols for reliability assessment; and integrating advanced privacy-preserving mechanisms into the adaptation process.
@@ 7 Conclusion
In this survey, we have presented a systematic analysis of LoRA, examining its theoretical underpinnings, technical advances, and diverse applications in adapting foundation models. The extensive adoption of LoRA across diverse domains—from natural language processing and computer vision to speech recognition and scientific computing—highlight its versatility and effectiveness. Its success in maintaining model performance while significantly reducing computational and storage requirements has made it particularly valuable for resource-constrained environments and specialized domain adaptations.
Despite these achievements, several critical challenges persist. The theoretical framework underlying LoRA’s effectiveness requires further development, particularly in understanding the interaction between low-rank adaptations and model capabilities. Additionally, questions regarding scalability, robustness, and secure deployment in production environments present ongoing research opportunities.

### v2

@@ 5.9.1 Language-vision Learning
Language-Vision Model-Based Adaptation. Sung et al. [229] adapted vision-language models (VLMs) by fine-tuning CLIP-BART for visual question answering and image captioning. Ji et al. [230] enhanced the adversarial robustness of VLMs through clustering-based LoRA for cross-modal retrieval. Zong et al. [231] proposed MoVA, which efficiently routes and fuses multiple vision foundation models such as CLIP [15], DINOv2 [166], and SAM [7] through a coarse-to-fine adapter mechanism. In addition, D-MoLE [232] employs a layer-wise expert allocator to assign LoRA experts automatically across layers, resolving architectural conflicts and routing instructions to facilitate knowledge sharing among experts, thereby improving language-vision performance.
@@ 5.9.2 Language-Audio Learning
In language-audio learning tasks, LoRA has emerged as a pivotal technique for addressing the modality gap between audio and text representations in foundation models. Applications primarily fall into two categories: speech recognition and audio content understanding and generation.
Speech Recognition. Fathullah et al. [241] introduced a direct audio-embedding method in which LoRA adapts self-attention layers for audio-text alignment, enabling large language models to process audio inputs effectively. Yusuf et al. [242] proposed speculative speech recognition by combining an RNN-Transducer-based ASR system [243] with an audio-prefixed language model adapted via LoRA, allowing the model to generate speculative predictions. Palaskar et al. [244] further introduced FLoRA with modality-specific adapters that can be selectively engaged, providing a robust framework for multimodal integration.
Speech Content Understanding and Generation. LOAE [245] combines CED-based audio encoding [246] with LLaMA-based text decoding for automated audio captioning, using a Q-Former [247] bridge and LoRA for efficient optimization. Qin et al. [211] introduced ATFLRec, a multimodal recommender system that leverages separate LoRA modules to fine-tune audio and text modalities in LLMs, demonstrating superior performance through a fusion-based approach.
@@ 6 Challenges and Discussion
While LoRA has excelled in efficiently adapting foundation models across diverse domains, several critical challenges and opportunities require further investigation.
Theoretical Understanding. Existing theoretical frameworks primarily focus on simplified settings or specific architectures [79], leaving gaps in our understanding of LoRA’s behavior in more complex scenarios. For instance, the interaction between LoRA adaptations and the pretrained model’s knowledge is not fully understood, particularly in how LoRA preserves useful features while modifying task-specific ones. Developing more comprehensive theoretical frameworks that can explain LoRA’s effectiveness across different architectures and tasks would be valuable for guiding future improvements.
@@
Computational Efficiency. The scalability of LoRA becomes increasingly crucial as LLMs continue to expand. Managing concurrent adaptations while handling variable-length sequences presents significant challenges in memory usage and computational overhead. The dynamic management of adapter modules and KV cache tensors can lead to memory fragmentation and increased I/O costs [148]. Advanced serving architectures and optimization techniques are essential for maintaining low latency in production environments, particularly when dealing with multiple concurrent adaptation requests.
Robustness and Verification. The deployment of LoRA-adapted models in critical applications necessitates robust verification methods. Current research inadequately addresses model behavior under distribution shifts and adversarial conditions [250]. Developing rigorous uncertainty quantification methods and formal verification techniques becomes paramount, especially for high-stakes applications in healthcare and autonomous systems, where model reliability directly impacts human safety.
@@
Future research directions should focus on establishing comprehensive theoretical frameworks that unify various LoRA design aspects; developing automated architecture search methods for optimal adapter configurations; creating efficient serving infrastructures for large-scale deployments; implementing standardized evaluation protocols for reliability assessment; and integrating advanced privacy-preserving mechanisms into the adaptation process.
@@ 7 Conclusion
In this survey, we have presented a systematic analysis of LoRA, examining its theoretical underpinnings, technical advances, and diverse applications in adapting foundation models. The extensive adoption of LoRA across diverse domains—from natural language processing and computer vision to speech recognition and scientific computing—highlights its versatility and effectiveness. Its success in maintaining model performance while significantly reducing computational and storage requirements has made it particularly valuable for resource-constrained environments and specialized domain adaptations. Despite these achievements, several critical challenges persist. The theoretical framework underlying LoRA’s effectiveness requires further development, particularly in understanding the interaction between low-rank adaptations and model capabilities. Additionally, questions regarding scalability, robustness, and secure deployment in production environments present ongoing research opportunities.